# Understanding Reader Perception Shifts upon Disclosure of AI Authorship

### v1

@@ 5.1. RQ1: Context is Critical: AI for Creation, Not Connection
Our results reveal a sharp distinction in how readers perceive AI authorship depending on the communicative purpose of the text. In acts of writing that are object-oriented, such as persuading (Convince), creating narratives (Imagine), or developing knowledge (Explore), AI involvement was viewed more favorably. The positive association with perceived competence in Convince and Imagine suggests that readers may appreciate AI as a tool that enhances the quality of argumentation or creative expression. This aligns with prior work suggesting that readers value AI’s contribution when it serves to consolidate or develop an author’s ideas (Hwang et al., 2025).
@@
In stark contrast, the use of AI in the person-oriented act of Interact was met with a strong negative reaction across nearly all perceptual measures. Our qualitative themes confirm that in such contexts, AI-authored text was perceived as “cold”, “impersonal” (Theme N1), and even “insincere” (Theme N4). This negative perception is consistent with prior work showing that AI disclosure can signal a lack of author engagement and thereby erode credibility (Jain et al., 2024; Jakesch et al., 2019; Schilke and Reimann, 2025). This finding highlights the importance of the ”communicative relationship” in interpersonal communication (Burleson, 2010). Because AI is perceived as incapable of genuine empathetic engagement (Liu et al., 2022; Rubin et al., 2025), its use in relational contexts can be interpreted as a violation of social expectations, making human attribution critical. Interestingly, in the Convince act, AI use boosted perceived competence without a corresponding penalty to likability or caring, suggesting that when AI’s role is to strengthen an argument, leveraging its perceived objectivity (Ovsyannikova et al., 2025), its use can be seen as strategic rather than socially detrimental.
5.2. RQ2: The Moderating Roles of AI Authorship Level and Reader Literacy
@@
Two factors emerged as key determinants of the magnitude of perception shifts: the disclosed level of AI contribution and the reader’s own AI literacy. As confirmed by our regression analysis, a higher disclosed AIRatio consistently led to more negative perceptions of the author. Our qualitative themes provide the reasoning behind this trend: as AI’s role grew, participants were more likely to perceive a loss of human touch (Theme N1), question the author’s expertise (Theme N3), and infer a lack of effort or agency (Theme N5). This finding aligns with recent work suggesting that dominant AI authorship can lower an author’s perceived credibility and ownership (Draxler et al., 2024; Kirk and Givi, 2025).
@@
Our results also confirm that a reader’s AI literacy can significantly mitigate these negative effects. Participants with higher self-reported AI literacy exhibited smaller negative perception shifts and, in some cases, expressed positive attitudes toward AI’s capabilities (Theme P8). This is consistent with prior studies indicating that greater familiarity with AI leads to a more pragmatic and less critical evaluation of its use (Li et al., 2024; Mahmud et al., 2024). AI literacy appears to function as a calibration mechanism, allowing readers to more accurately assess the role and utility of AI in the writing process. However, the effect size of AI literacy was smaller than that of the writing act itself. This implies that while literacy can soften negative reactions, it does not erase the strong social norms governing contexts like interpersonal communication, where even AI-literate readers may expect human-authored content.
5.3. RQ3: “Considerate” AI Authorship Can Be Socially Acceptable
@@
While the overall trend in our study was toward negative perception shifts, a substantial portion of responses (38.6%) indicated neutral or even positive changes. These reactions were not random; they were most common when the disclosed AI contribution was low (under 50%) and occurred in object-oriented acts such as Explore, Convince, and Imagine, rather than person-oriented acts like Interact.
@@
Our qualitative data reveal that this “considerate” AI authorship was perceived positively when it served a clear, assistive role that did not supplant the author’s voice. Participants noted that when AI was used for supplementary tasks, it could preserve the author’s core intention or even strengthen their argument (Theme P6). This aligns with findings from Hwang et al. (2025), who observed that writers themselves feel their work is authentic when AI is used to refine or develop their own pre-existing ideas, particularly in non-emotional writing. Therefore, our findings suggest that the social acceptability of AI authorship hinges on its perceived role: when positioned as a supportive tool that enhances human expression, it can be valued; when it is seen as a substitute for human effort and emotion, it is often rejected.
@@ 5.4. Design Implications for AI-mediated Writing
5.4.1. Design for Context-Sensitive Transparency
Our results show that a simple quantitative disclosure, such as stating the amount of AI contribution, is a blunt instrument that is often insufficient, particularly in relational contexts. Instead, transparency mechanisms should be context-sensitive and qualitative. Rather than focusing on how much AI was used, systems should help authors explain the nature and purpose of the AI’s involvement (e.g., AI assisted with grammatical revisions or AI generated the initial draft). This aligns with emerging best practices, such as the ACM’s policy on authorship, which calls for specifying the role AI played. Furthermore, the mode of disclosure should adapt to the communicative act. For descriptive or exploratory writing, explicitly showing the history of AI edits, such as with text highlighting (Hoque et al., 2024), may help to increase the credibility of authors. For creative co-writing, disclosure could emphasize the collaborative nature of the process. In sensitive interpersonal contexts, transparency should focus on preserving the human author’s effort and emotional investment.
@@ 5.4.2. Preserve and Communicate Human Effort
Our qualitative analysis reveals that negative reactions to AI disclosure are significantly mitigated when readers can perceive authentic human effort. This suggests that readers evaluate not just the extent of AI authorship, but the degree of perceptible human agency in the final text. Therefore, AI writing tools should be designed not only to disclose AI’s role but also to actively preserve and communicate the author’s contributions. Interfaces could integrate features that visualize human effort, such as edit histories, authorship heatmaps, or representations of the author’s strategic decisions and intentions (Peng et al., 2025; Wang et al., 2025b). By foregrounding the author’s role in guiding, refining, and curating the text, systems can frame human-AI collaboration as a process of thoughtful engagement rather than delegation. This approach complements context-sensitive transparency: while the latter aims to minimize harm from disclosure, making human effort visible can actively restore authenticity and trust.
@@ 5.4.3. Provide Adaptive Support in Emotional and Social Contexts
The strongest negative perception shifts occurred in writing with emotional or relational significance, where readers expect sincerity and personal investment. This finding underscores that the social function of writing fundamentally shapes how AI authorship is judged. AI-assisted writing systems should therefore be designed with social and emotional awareness. Interfaces could be developed to detect communicative intent, for example by recognizing linguistic markers of empathy, apology, or gratitude. In such contexts, the system could adapt its support, perhaps by limiting its suggestions to structural or stylistic refinements while leaving core affective expressions to the human author. It could also prompt writers to add a personal reflection or review the text to ensure it conveys their genuine feelings. Such adaptive designs would not only help preserve relational authenticity but also guide authors toward a more balanced and appropriate division of labor with AI.
@@ 5.4.4. Foster AI Literacy for Both Authors and Readers
Our study confirms that higher AI literacy correlates with a greater tolerance for AI authorship. This suggests that perceptions are shaped not only by the act of writing but also by the user’s understanding of AI’s role. We therefore see enhancing AI literacy as an important step toward fostering calibrated trust. AI writing tools can be designed to actively cultivate this literacy. For authors, reflective interfaces could visualize their reliance patterns, such as the frequency with which they accept suggestions verbatim versus editing them, and prompt them to consider their authorship balance. For readers, provenance interfaces could provide contextual information about typical forms of AI assistance, helping them to calibrate their trust without overreacting to the mere fact of disclosure. Over time, such feedback loops could help establish shared social norms around what constitutes appropriate and considerate use of AI in writing.
@@ 5.5. Limitations and Future Work
First, we manipulated the perceived level of AI authorship using a quantitative ratio of sentences, which, while systematic, does not account for the semantic importance of those sentences. A sentence containing the main thesis of an argument carries more weight than one providing a minor detail; consequently, the perceived influence of AI likely varied depending on which specific sentences were highlighted. Future work should explore how readers’ perceptions are affected when AI contributes to different functional parts of a text, such as core arguments versus stylistic refinements.
Second, our simulation of AI disclosure does not capture the full complexity of real-world human-AI writing collaborations. As suggested by prior work, a more nuanced representation would involve describing the nature of the collaboration, such as the types of prompts the author used, the kinds of assistance the AI provided, and how the author accepted, rejected, or modified the AI’s suggestions (Draxler et al., 2024; He et al., 2025). Future experiments should include these contribution-level descriptions to better model the diverse patterns of human-AI co-authorship.
Finally, our study was conducted exclusively with Japanese participants. While our design was not language-specific, cultural norms surrounding communication, authenticity, and technology may influence how AI authorship is perceived. The generalizability of our findings to other linguistic and cultural contexts is therefore an open question. Cross-cultural comparative studies are needed to investigate how these nuanced social dynamics vary across different populations.
@@ 6. Conclusion
We examined how the disclosure of AI authorship in writing influences readers’ perceptions across different acts of writing. Our user study with 261 participants revealed that disclosure generally erodes trust, caring, competence, and likability, particularly in social and emotional contexts. However, we found that these negative effects are moderated by a reader’s AI literacy and by the extent to which human effort and agency remain perceptible. Our findings demonstrate that the social acceptance of AI-mediated writing depends not just on the content but on how human intentionality is communicated and perceived. Based on these insights, we offer design implications for AI writing tools that emphasize context-sensitive transparency, visibility of human effort, and literacy building to sustain authenticity. Our work calls for re-imagining AI writing assistance as a co-creative partnership and relational process where technology amplifies, rather than replaces, human expression and empathy. Future research should explore these dynamics across different cultures and over time as AI becomes more deeply integrated into our communicative practices.
@@ 7. GenAI Usage Disclosure
Finally, throughout the writing process, we used Gemini, GPT-5, and Writefull to improve the clarity, style, and grammatical correctness of the manuscript.
Acknowledgement
@@ Appendix A Statements in the Author Evaluation and Post-experimental Questionnare
Table 5. Author’s impression evaluation statements used on pre/post-disclosure:
@@
Table 6. AI literacy assessment used on post-experiment questionnaire:
@@ Appendix B Detailed Result of Regression Model
Table 7. Mixed Linear Model Regression Results for TrustworthinessShift. Marginal R²: 0.2173, Conditional R²: 0.4512:
@@
Table 8. Mixed Linear Model Regression Results for CaringShift. Marginal R²: 0.2104, Conditional R²: 0.3440:
@@
     To Convince -1.058 0.612 -1.730 0.084 -2.256 0.141
     To Explore 1.652 0.614 2.692 0.007** 0.449 2.855
     To Imagine 0.465 0.616 0.755 0.450 -0.743 1.673
     To Interact -2.984 0.617 -4.838 0.000*** -4.192 -1.775
     To Reflect -0.048 0.617 -0.077 0.939 -1.257 1.162
AI Ratio -2.874 0.451 -6.373 0.000*** -3.757 -1.990
Apply AI Literacy 0.406 0.251 1.619 0.105 -0.086 0.898
Detect AI Literacy 0.335 0.293 1.142 0.253 -0.240 0.910
AI Persuasion Literacy 0.048 0.277 0.173 0.862 -0.495 0.590
@@
     AI Ratio : To Convince -0.383 0.639 -0.599 0.549 -1.636 0.870
     AI Ratio : To Explore 0.905 0.641 1.410 0.158 -0.353 2.162
     AI Ratio : To Imagine 1.400 0.641 2.183 0.029* 0.143 2.657
     AI Ratio : To Interact -0.482 0.638 -0.756 0.450 -1.733 0.769
     AI Ratio : To Reflect 1.207 0.644 1.876 0.061 -0.054 2.468
AI Ratio: Apply AI Literacy 0.337 0.185 1.821 0.069 -0.026 0.700
AI Ratio: AI Persuasion Literacy 0.435 0.183 2.375 0.018* 0.076 0.794
Group Var 6.006 0.280
@@
Table 9. Mixed Linear Model Regression Results for CompetenceShift. Marginal R²: 0.1385, Conditional R²: 0.3337:
@@
Table 10. Mixed Linear Model Regression Results for LikabilityShift. Marginal R²: 0.2227, Conditional R²: 0.2920:
@@
Table 11. Mixed Linear Model Regression Results for FutureShift. Marginal R²: 0.1435, Conditional R²: 0.2245:

### v2

@@ 5.1. RQ1: Context is Critical: AI for Creation, Not Connection
Our results reveal a sharp distinction in how readers perceive AI authorship depending on the communicative purpose of the text (Table 3). In acts of writing that are object-oriented, such as persuading (Convince), creating narratives (Imagine), or developing knowledge (Explore), AI involvement was viewed more favorably. This task-dependent favoritism aligns with insights from Castelo et al. (2019), who found that objective tasks, such as predicting stock market outcomes, were more readily accepted than subjective tasks like romantic partner recommendation. However, AI generation may still be discouraged in objective tasks where accuracy is paramount, particularly in forecasting contexts (Dietvorst et al., 2015). The positive association with perceived competence in Convince and Imagine suggests that readers may appreciate AI as a tool that enhances the quality of argumentation or creative expression. This is supported by prior work suggesting that readers value AI’s contribution when it serves to consolidate or develop an author’s ideas (Hwang et al., 2025).
@@
In stark contrast, the use of AI in the person-oriented act of Interact elicited strong negative reactions. Our qualitative themes confirm that in such contexts, AI-authored text was perceived as “cold”, “impersonal” (Theme N1), and even “insincere” (Theme N4). This negative perception is consistent with prior work showing that AI disclosure can signal a lack of author engagement and thereby erode credibility (Jain et al., 2024; Jakesch et al., 2019; Schilke and Reimann, 2025). These findings underscore the context sensitivity of algorithm aversion. As noted by Morewedge (2022), algorithm aversion is often triggered by tasks lacking established evaluation criteria (e.g., creative work) or those dependent on individual subjectivity (e.g., gift recommendations). The Interact act embodies both characteristics, and our findings confirm it to be the act of writing most sensitive to AI disclosure. This also highlights the importance of the “communicative relationship” in interpersonal communication (Burleson, 2010). Because AI is perceived as incapable of genuine empathetic engagement (Liu et al., 2022; Rubin et al., 2025), its use in relational contexts can be interpreted as a violation of social expectations, making human attribution critical. Interestingly, in the Convince and Interact act, AI use boosted perceived competence, suggesting that when AI’s role is to strengthen an argument and leverage its perceived objectivity (Ovsyannikova et al., 2025), its use can be seen as strategic rather than socially detrimental.
5.2. RQ2: Moderating Roles of AI Authorship Level and Readers’ Literacy
@@
Two factors emerged as key determinants of the magnitude of perception shifts: the disclosed level of AI contribution and the reader’s AI literacy. As confirmed by our regression analysis, a higher disclosed AIRatio consistently led to more negative perceptions of the author. Our qualitative themes provide the reasoning behind this trend: as AI’s role grew, participants were more likely to perceive a loss of human touch (Theme N1), question the author’s expertise (Theme N3), and infer a lack of effort or agency (Theme N5). This echoes recent work suggesting that dominant AI authorship can lower an author’s perceived credibility and ownership (Draxler et al., 2024; Kirk and Givi, 2025).
@@
Our results also confirm that a reader’s AI literacy can significantly mitigate these negative effects. Participants with higher self-reported AI literacy exhibited smaller negative perception shifts and, in some cases, expressed positive attitudes toward AI’s capabilities (Theme P8). This is consistent with prior studies indicating that greater familiarity with AI and positive past interactions with AI foster a more pragmatic and less critical evaluation of its use (Böhm et al., 2023; Li et al., 2024b; Mahmud et al., 2024). AI literacy appears to function as a calibration mechanism, allowing readers to more accurately assess the role and utility of AI in the writing process. However, the effect size of AI literacy was smaller than that of the writing act itself. This implies that while literacy can soften negative reactions, it does not erase the strong social norms governing contexts like interpersonal communication, where even AI-literate readers often expect human-authored content.
5.3. RQ3: Assistive AI Authorship Can Be Socially Acceptable
@@
While the overall trend in our study was toward negative perception shifts, a substantial portion of responses (38.6%) indicated neutral or even positive changes. These reactions were most common when the disclosed AI contribution was low (under 50%) and occurred in descriptive or argumentative acts such as Explore, Convince, and Imagine, rather than interpersonal acts like Interact.
@@
Our qualitative data revealed that such assistive AI authorship was perceived positively when it served a clear, assistive role that did not supplant the author’s voice. This preference may stem from human favoritism, where human authorship serves as a cue for trust (Buchanan and Hickman, 2024; Zhang and Gosline, 2023). Participants agreed that when AI was used for supplementary tasks, the author’s core intention and argument can be preserved, or even strengthened (Theme P6). This aligns with prior findings that limiting a machine’s capability to an advisory role can lead to greater permissibility in decision-making processes (Bigman and Gray, 2018). In the context of non-emotional writing, Hwang et al. (2025) similarly found that writers perceive their work as authentic when AI is used to refine or develop their own pre-existing ideas. Our findings confirm that the social acceptability of AI authorship hinges on its perceived role: when positioned as a supportive tool that enhances human expression, it can be valued; when it is seen as a substitute for human effort and emotion, it is often rejected.
@@ 5.4. Design Implications for AI-mediated Writing
5.4.1. Design for Functional, Context-Sensitive Transparency
Our results showed that simple quantitative disclosure, such as stating the amount of AI contribution, was a blunt and insufficient instrument, particularly in relational contexts. Moreover, recent research suggests that readers often anticipate others will use AI tools more than they themselves (Purcell et al., 2024). This social assumption may raise suspicion regarding AI involvement, potentially eroding trust in the author (Jakesch et al., 2019). Consequently, we believe that transparency mechanisms should disclose the functional use of AI in a context-sensitive manner. For example, rather than focusing on how much AI was used, systems should help authors explain the nature and purpose of AI involvement (e.g., “AI assisted with grammatical revisions” or “AI generated the initial draft”). This aligns with emerging best practices, such as the ACM’s policy on authorship, which calls for specifying the role AI played (Association for Computing Machinery (ACM), 2025).
This functional transparency is particularly relevant for non-native speakers, who often use AI to ensure linguistic accuracy rather than to offload cognitive effort (Hwang et al., 2023; Li et al., 2024a). In such cases, explicitly disclosing that AI was used to overcome language barriers can frame the assistance as a pragmatic accessibility aid. This context may induce reader sympathy and understanding, thereby mitigating the negative impact of disclosure by highlighting the author’s effort to communicate effectively in a non-native language.
Furthermore, the mode of disclosure should adapt to the communicative act. For descriptive or exploratory writing, explicitly showing the history of AI edits, such as with text highlighting (Hoque et al., 2024), may help increase author credibility. For creative co-writing, disclosure could emphasize the collaborative nature of the process. In sensitive interpersonal contexts, transparency should focus on preserving the human author’s effort and emotional investment.
@@ 5.4.2. Preserve and Communicate Human Effort
Our qualitative analysis revealed that negative reactions to AI disclosure were significantly mitigated when authentic human effort remained perceptible. This suggests that readers evaluate not just the extent of AI authorship, but the degree of perceptible human agency in the final text. Therefore, AI writing tools should be designed not only to disclose the role of AI but also to actively preserve and communicate the author’s contributions. Interfaces could integrate features that visualize human effort, such as edit histories, authorship heatmaps, or representations of the author’s strategic decisions and intentions (Peng et al., 2025; Wang et al., 2025b). By foregrounding the author’s agency (e.g., the acts of guiding, initiating, and curating the text as conceptualized by Rae (2024)), systems can frame human-AI collaboration as a process of thoughtful engagement rather than delegation. This approach complements context-sensitive transparency. As social and emotional contexts place a premium on human effort, systems should explicitly and visibly convey human touch to proactively restore authenticity and trust.
@@ 5.4.3. Provide Adaptive Support in Emotional and Social Contexts
The strongest negative perception shifts occurred in writing with emotional or relational significance, where readers expect sincerity and personal investment. This underscores that the social function of writing fundamentally shapes how AI authorship is judged. AI-assisted writing systems may therefore benefit from being designed with social and emotional awareness. For example, they could detect communicative intent by recognizing linguistic markers of empathy, apology, or gratitude and alert users to the social risks of automation and suggest drafting the content personally. Alternatively, it could adapt its support by limiting suggestions to structural or stylistic refinements while leaving core affective expressions to the human author. Systems could also prompt writers to add personal reflections or review the text to ensure it conveys genuine feelings. Such adaptive designs would not only help preserve relational authenticity but also guide authors toward a more balanced and appropriate division of labor with AI.
@@ 5.4.4. Foster AI Literacy for Both Authors and Readers
Our study confirms that higher AI literacy correlates with greater tolerance for AI authorship. This suggests that perceptions are shaped not only by the act of writing but also by the user’s understanding of the role of AI. We therefore view enhancing AI literacy as an important step toward fostering calibrated trust. AI writing tools can be designed to actively cultivate this literacy. For authors, reflective interfaces could visualize their reliance patterns, such as the frequency with which they accept suggestions verbatim versus editing them, and prompt them to consider their authorship balance. For readers, provenance interfaces could provide contextual information about typical forms of AI assistance, helping them calibrate trust without overreacting to the mere fact of AI disclosure, thereby allowing them to better perceive the author’s dedication. Over time, such feedback loops could help establish shared social norms around what constitutes appropriate and considerate use of AI in writing.
@@ 5.5. Limitations and Future Work
First, we manipulated the perceived level of AI authorship using a quantitative ratio of sentences. While systematic, this does not account for the semantic weight of those sentences. For instance, a sentence articulating a core thesis carries more weight than one providing a minor detail; consequently, the perceived influence of AI likely varied depending on which specific sentences were highlighted. Future work should explore how reader perceptions are affected when AI contributes to different functional parts of a text, such as core arguments versus stylistic refinements.
Second, our simulation of AI disclosure simplifies the complexity of real-world human-AI writing collaborations. As suggested by prior work, a more nuanced representation would involve describing the nature of the collaboration, such as the types of prompts used, the specific assistance provided by the AI, and how the author accepted, rejected, or modified those suggestions (Draxler et al., 2024; He et al., 2025; Rae, 2024). Future experiments should include these contribution-level descriptions to better model diverse patterns of human-AI co-authorship.
Third, our study was conducted exclusively with Japanese participants. While our design was not language-specific, cultural norms surrounding communication, authenticity, and technology may influence how AI authorship is perceived. The generalizability of our findings to other linguistic and cultural contexts remains an open question, and cross-cultural comparative studies are needed to investigate how these nuanced social dynamics vary across different populations.
Finally, we assessed AI literacy as a static personal attribute. However, AI literacy and attitudes toward AI are likely to evolve over time. Future longitudinal research is necessary to determine whether our findings persist as AI technologies and user familiarity continue to mature.
@@ 6. Conclusion
We examined how the disclosure of AI authorship in writing influences readers’ perceptions across different acts of writing. Our study with 261 participants revealed that disclosure generally erodes perceptions of trustworthiness, caring, competence, and likability, particularly within social and emotional contexts. However, these negative effects are moderated by the reader’s AI literacy and the degree to which human effort and agency remain visible. Our findings demonstrate that the social acceptability of AI-mediated writing hinges not merely on textual quality, but on how human intentionality is communicated and perceived. Based on these insights, we propose design implications for AI writing tools that prioritize context-sensitive transparency, the preservation of human effort, and the cultivation of AI literacy to sustain authenticity. Our work advocates for re-imagining AI writing assistance as a co-creative partnership and relational process where technology amplifies, rather than replaces, human expression and empathy. As AI becomes increasingly embedded in our communicative practices, future research must continue to explore these dynamics across diverse cultural contexts and longitudinal timelines.
@@ 7. GenAI Usage Disclosure
Finally, throughout the writing process of this manuscript, we used Gemini, GPT-5, and Writefull to improve the clarity, style, and grammatical correctness of the manuscript.
Acknowledgements.
@@ Appendix A Statements in the Author Evaluation and Post-experimental Questionnare
Table 5. Author’s impression evaluation statements used on pre/post-disclosure
Author’s impression evaluation statements used on pre/post-disclosure
@@
Table 6. AI literacy assessment used on post-experiment questionnaire
Participant’s AI Literacy Assessment used on post-experiment questionnaire
@@ Appendix B Detailed Result of Regression Model
Table 7. Mixed Linear Model Regression Results for TrustworthinessShift. Marginal R²: 0.2173, Conditional R²: 0.4512
Mixed linear model regression table for TrustworthinessShift with Act type (6 categories, reference: To Describe), AIRatio, and AI literacy measures as predictors. Shows coefficients, standard errors, z-values, p-values, and 95% confidence intervals. Significant main effects include AIRatio (negative) and To Interact act type (negative). Significant positive interaction between AIRatio and Apply AI Literacy.
@@
Table 8. Mixed Linear Model Regression Results for CaringShift. Marginal R²: 2.036, Conditional R²: 0.3152
@@
     To Convince -1.075 0.619 -1.737 0.082 -2.287 0.138
     To Explore 1.612 0.620 2.601 0.009** 0.397 2.826
     To Imagine 0.410 0.622 0.659 0.510 -0.809 1.628
     To Interact -3.050 0.622 -4.903 0.000*** -4.269 -1.831
     To Reflect -0.144 0.622 -0.232 0.817 -1.363 1.075
AIRatio -2.838 0.454 -6.251 0.000*** -3.727 -1.948
Detect AI Literacy 0.511 0.265 1.927 0.054 -0.009 1.031
AI Persuasion Literacy -0.001 0.266 -0.003 0.997 -0.522 0.521
@@
     AIRatio : To Convince -0.390 0.643 -0.606 0.544 -1.651 0.871
     AIRatio : To Explore 0.892 0.645 1.383 0.167 -0.372 2.157
     AIRatio : To Imagine 1.372 0.644 2.129 0.033* 0.109 2.635
     AIRatio : To Interact -0.508 0.641 -0.792 0.428 -1.765 0.749
     AIRatio : To Reflect 1.098 0.646 1.700 0.089 -0.168 2.365
AIRatio: AI Persuasion Literacy 0.467 0.184 2.538 0.011* 0.106 0.828
Group Var 4.951 0.241
@@
Table 9. Mixed Linear Model Regression Results for CompetenceShift. Marginal R²: 0.1385, Conditional R²: 0.3337
Mixed linear model regression table for CompetenceShift with Act type (6 categories, reference: To Describe), AIRatio, and AI literacy measures as predictors. Shows coefficients, standard errors, z-values, p-values, and 95% confidence intervals. Significant main effects include AIRatio (negative), To Convince act type (positive), To Imagine act type (positive), and To Interact act type (positive). Significant positive interaction between AIRatio and Apply AI Literacy.
@@
Table 10. Mixed Linear Model Regression Results for LikabilityShift. Marginal R²: 0.2227, Conditional R²: 0.2920
Mixed linear model regression table for LikabilityShift with Act type (6 categories, reference: To Describe), AIRatio, and AI literacy measures as predictors. Shows coefficients, standard errors, z-values, p-values, and 95% confidence intervals. Significant main effects include AIRatio (negative), To Interact act type (negative), and Detect AI Literacy (positive). Significant positive interaction between AIRatio and Detect AI Literacy.
@@
Table 11. Mixed Linear Model Regression Results for FutureShift. Marginal R²: 0.1435, Conditional R²: 0.2245
Mixed linear model regression table for FutureShift with AIRatio and AI literacy measures as predictors. Shows coefficients, standard errors, z-values, p-values, and 95% confidence intervals. Significant main effect of AIRatio (negative). Significant positive interaction between AIRatio and AI Persuasion Literacy.